Show a complete solution.
Big plane. A six-passenger plane cruises at 180 mph in
calm air. If the plane flies 7 miles with the wind in the
same amount of time as it flies 5 miles against the wind,
then what is the wind speed?

Answer by checkley77(12844)

D=RT OR T=D/R
T=7/(180+W)
T=5/(180-W) SEEING AS THE TIMES ARE EQUAL WE SET THESE 2 EQUATION EQUAL
7/(180+W)=5/(180-W)
7(180-W)=5(180+W)
1,260-7W=900+5W
-7W-5W=900-1,260
-12W=-360
W=-360/-12
W=30 MPH.IS THE WIND SPEED.
PROOF:
7/(180+30)=5/(180-30)
7/210=5/150
1/30=1/30